# Position: Bayesian Deep Learning is Needed in the Age of Large-Scale AI

### v3

@@ 2.2 Data Efficiency
Unlike many machine learning approaches that may require large datasets to generalize effectively, BDL leverages prior knowledge and updates beliefs as new data become available. This allows BDL to extract meaningful information from small datasets, making it more efficient in scenarios where collecting large amounts of data is challenging or costly (Finzi et al., 2021; Immer et al., 2022; Shwartz-Ziv et al., 2022; Schwöbel et al., 2022; van der Ouderaa et al., 2023). In addition, the regularization effect introduced by the probabilistic nature of its Bayesian approach is beneficial in preventing overfitting and contributing to better generalization from fewer samples (Rothfuss et al., 2022; Sharma et al., 2023). BDL’s uncertainty modeling helps resist the influence of outliers, making it well-suited for real-world scenarios with noisy or out-of-distribution data. This also makes it attractive for foundation model fine-tuning, where data are commonly small and sparse, and uncertainty is important.
@@ 3.3 Posterior Sampling Algorithms
Furthermore, SG-MCMC is still considered expensive for deep learning applications. A step forward in this regard would be to learn from the machine learning and systems community how to make Monte Carlo faster using contemporary hardware (Zhang et al., 2022a; Wang et al., 2023b). Algorithms such as Stein variational gradient descent (SVGD; Liu & Wang, 2016) occupy a middle ground between optimization and sampling, by employing optimization-type updates but with a set of interacting particles. While recent advances show promising results in BNN settings (D’Angelo et al., 2021; D’Angelo & Fortuin, 2021; Pielok et al., 2022), these methods often perform poorly in high-dimensional problems. Alternatively, convergence rates and posterior exploration can be improved with cyclical step-size schedules (Zhang et al., 2020).
@@ A.2 Variational Inference
with respect to a variational distribution $q(\boldsymbol{\theta})$ within some variational family of distributions $\mathcal{Q}_{\boldsymbol{\Theta}}$ (Wainwright & Jordan, 2008). In (3), $\mathbb{D}_{\mathrm{KL}}$ denotes the Kullback-Leibler (KL) divergence. Since the posterior $p(\boldsymbol{\theta}\mid\mathcal{D})$ is the distribution to be approximated and as such is not accessible, the variational problem described by (3) cannot be solved directly. However, it can be shown that solving this variational problem is mathematically equivalent to maximizing the variational objective
@@
In general, variational inference is not guaranteed to converge to the posterior $p(\boldsymbol{\theta}\mid\mathcal{D})$ unless the variational objective is convex in the variational parameters and the posterior is a member of the variational family, that is, $p(\boldsymbol{\theta}\mid\mathcal{D})\in\mathcal{Q}_{\boldsymbol{\Theta}}$. Various approximate inference methods have been developed to solve the variational problem described by (3). These methods make different assumptions about the variational family $\mathcal{Q}_{\boldsymbol{\Theta}}$, and therefore result in different posterior approximations.
@@ A.3 Ensembles
Deep ensembling refers to a procedure where a neural network architecture is re-trained multiple times with different initializations to find different parameter settings, and then the resulting predictive distributions at those parameter settings are averaged at test time (Lakshminarayanan et al., 2017). In practice, deep ensembles provide a simple approach to representing epistemic uncertainty by capturing the variability in model predictions. This approach contrasts with more conventional Bayesian methods that involve sampling from or optimizing the posterior distribution. Although unorthodox as an approximate inference procedure, deep ensembles often provide a closer approximation to the true posterior predictive distribution than many conventional approximate inference methods in deep learning (Wilson & Izmailov, 2020; Izmailov et al., 2021b), such as variational inference with a Gaussian approximate posterior.
@@
There are different ways to interpret this predictive distribution. Many works have explored the connections between Bayesian inference and deep ensembles, such as those by Wilson & Izmailov (2020); Gustafsson et al. (2020); He et al. (2020); Pearce et al. (2020); D' Angelo & Fortuin (2021). One interpretation views Equation 4 as a Monte Carlo approximation of Equation 5, where the posterior of the parameters is represented as a set of point masses centered at different modes, which may be viewed as approximate posterior samples. However, this interpretation is not the most insightful. Instead, approximate inference can be viewed as the task of accurately approximating the integral in Equation 5. From this perspective, the focus is not on collecting posterior samples. For a fixed computational budget, a Monte Carlo average of predictive distribution values based on exact posterior samples can provide a poor approximation of the integral relative to alternatives. A more compelling approach to numerical integration is to choose parameter values that represent typical points in the posterior, indicative of regions with significant posterior probability mass, and that yield diverse predictions on the test set. A natural heuristic to achieve this goal is to choose points corresponding to different posterior modes, as achieved by deep ensembles (Wilson & Izmailov, 2020). In practice, there is more functional variability across different posterior modes compared to samples in the vicinity of a single mode, such as the ones found from a variational Gaussian approximation of the posterior.
@@ A.4 Posterior Sampling Algorithms
Simulating from the following continuous-time stochastic differential equation (SDE) produces samples with $p(\boldsymbol{\theta}\mid\mathcal{D})$ as its stationary distribution:
@@
where $\alpha>0$ is the step size parameter and $\boldsymbol{\xi}$ is standard Gaussian noise. This discrete-time algorithm is known as the unadjusted Langevin algorithm, or the Langevin Monte Carlo algorithm. However, unlike the continuous-time Langevin diffusion equation, the discrete-time unadjusted Langevin algorithm does not simulate samples with $p(\boldsymbol{\theta}\mid\mathcal{D})$ as its stationary distribution, but instead produces samples that are approximately drawn from $p(\boldsymbol{\theta}\mid\mathcal{D})$. The discretization of the SDE leads to a bias in the posterior samples, which can be reduced by decreasing the step size parameter $\alpha$.
@@
Other notable variants of SG-MCMC include stochastic gradient HMC (SG-HMC; Chen et al., 2014), which accelerates convergence using auxiliary momentum variables, and cyclical SG-MCMC (Zhang et al., 2020), which employs a cyclical step size schedule to efficiently explore multiple modes of the parameter posterior distribution.
@@ A.5 Prior Specification
which corresponds to $l_{2}$ regularization in the sense of maximum a posteriori (MAP) solutions with $\lambda=1/(2\sigma^{2})$. Thus, the larger the prior variance, the less regularization is incorporated, and vice versa. Combined with specific activation functions, such as the logistic function, which is close to linear around 0, choosing a small $\sigma^{2}$ results in more linear behavior of the neurons and their compositions, while large $\sigma^{2}$ allows for more non-linear behavior. Thus, the popular approach of choosing standard Gaussian priors is not satisfactory in most cases and may lead to misspecified models. This, in turn, can cause the cold posterior effect that has been known to be the case for linear models (Grünwald & Van Ommen, 2017), but remains problematic for BNNs (Wenzel et al., 2020; Fortuin et al., 2022). For a specific problem, $\sigma^{2}$ can be chosen via hyper-parameter tuning or empirical Bayes. Moreover, a direct translation of the tuned $\sigma^{2}$ for some architecture (or, equivalently, $\lambda$ for frequentist neural networks) is possible. Another approach is to impose an inverse-Gamma hyper-prior on $\sigma^{2}$, for example, $\sigma^{2}\sim\Gamma^{-1}(\alpha,\beta)$, see Lampinen & Vehtari (2001). To incorporate prior dependencies between the parameters, i.i.d. Gaussian priors can be extended to multivariate normals with a zero mean vector and a covariance matrix $\boldsymbol{\Sigma}$, i.e.,
@@ A.7 Deep Kernel Machines
The objective in (10) is maximized when $\mathbf{G}_{j}=\mathbf{K}(\mathbf{G}_{j-1})$, which is the same as the corresponding NNGP (Lee et al., 2017; Agrawal et al., 2020). If tempering is carried out according to the width with $\beta_{j}=1/r$, then the following objective is obtained:
@@
A model that optimizes the objective in (11) is called a DKM, and (11) is known as the DKM objective. In the limit, the DKM objective does not depend on intermediate features $\mathbf{F}_{j}$, which means that learned representations in a DKM are described entirely by deterministic Gram matrices $\mathbf{G}_{1},\ldots,\mathbf{G}_{\eta}$. To interpret the DKM objective, notice that the likelihood term encourages data fitting, and the KL divergences regularize the model toward the NNGP (Lee et al., 2017; Agrawal et al., 2020). The amount of representation learning in the DKM can be controlled by varying the $\rho_{j}$ parameters. In contrast, the lack of likelihood term in the NNGP objective (10) prevents representation learning from occurring in the NNGP; intermediate Gram matrices are fixed and depend only on the input data.

### v5

@@ 2.2 Data Efficiency
Unlike many machine learning approaches that may require large datasets to generalize effectively, BDL leverages prior knowledge and updates beliefs as new data become available. This allows BDL to extract meaningful information from small datasets, making it more efficient in scenarios where collecting large amounts of data is challenging or costly (Finzi et al., 2021; Immer et al., 2022b; Shwartz-Ziv et al., 2022; Schwöbel et al., 2022; van der Ouderaa et al., 2023). In addition, the regularization effect introduced by the probabilistic nature of its Bayesian approach is beneficial in preventing overfitting and contributing to better generalization from fewer samples (Rothfuss et al., 2022; Sharma et al., 2023). BDL’s uncertainty modeling helps resist the influence of outliers, making it well-suited for real-world scenarios with noisy or out-of-distribution data. This also makes it attractive for foundation model fine-tuning, where data are commonly small and sparse, and uncertainty is important.
@@ 3.3 Posterior Sampling Algorithms
Furthermore, SG-MCMC is still considered expensive for deep learning applications. A step forward in this regard would be to learn from the machine learning and systems community how to make Monte Carlo faster using contemporary hardware (Zhang et al., 2022a; Wang et al., 2023b). Algorithms such as Stein variational gradient descent (SVGD; Liu & Wang, 2016) occupy a middle ground between optimization and sampling, by employing optimization-type updates but with a set of interacting particles. While recent advances show promising results in BNN settings (D’Angelo et al., 2021; D’Angelo & Fortuin, 2021; Pielok et al., 2022), these methods often perform poorly in high-dimensional problems. Alternatively, convergence rates and posterior exploration can be improved with cyclical step-size schedules (Zhang et al., 2020b).
@@ A.1 Laplace Approximations
An advantage of the Laplace approximation is that it enables to compute the marginal likelihood of the approximate posterior in closed form (Immer et al., 2021a), which can be used for Bayesian model selection in neural networks, for instance, for invariance learning (Immer et al., 2022b), linguistic probing of language models (Immer et al., 2022a), or neural network pruning (Dhahri et al., 2024). Thus, the Laplace approximation makes BDL more computationally feasible.
@@ A.2 Variational Inference
with respect to a variational distribution $q(\boldsymbol{\theta})$ within some variational family of distributions $\mathcal{Q}_{\boldsymbol{\Theta}}$ (Wainwright & Jordan, 2008). In expression (3), $\mathbb{D}_{\mathrm{KL}}$ denotes the Kullback-Leibler (KL) divergence. Since the posterior $p(\boldsymbol{\theta}\mid\mathcal{D})$ is the distribution to be approximated and as such is not accessible, the variational problem described by expression (3) cannot be solved directly. However, it can be shown that solving this variational problem is mathematically equivalent to maximizing the variational objective
@@
In general, variational inference is not guaranteed to converge to the posterior $p(\boldsymbol{\theta}\mid\mathcal{D})$ unless the variational objective is convex in the variational parameters and the posterior is a member of the variational family, that is, $p(\boldsymbol{\theta}\mid\mathcal{D})\in\mathcal{Q}_{\boldsymbol{\Theta}}$. Various approximate inference methods have been developed to solve the variational problem described by expression (3). These methods make different assumptions about the variational family $\mathcal{Q}_{\boldsymbol{\Theta}}$, and therefore result in different posterior approximations.
@@ A.3 Ensembles
Deep ensembling refers to a procedure where a neural network architecture is re-trained multiple times with different initializations to find different parameter settings, and then the resulting predictive distributions at those parameter settings are averaged at test time (Lakshminarayanan et al., 2017). In practice, deep ensembles provide a simple approach to representing epistemic uncertainty by capturing the variability in model predictions. This approach contrasts with more conventional Bayesian methods that involve sampling from the posterior distribution. Although unorthodox as an approximate inference procedure, deep ensembles often provide a closer approximation to the true posterior predictive distribution than many conventional approximate inference methods in deep learning (Wilson & Izmailov, 2020; Izmailov et al., 2021b), such as variational inference with a Gaussian approximate posterior.
@@
There are different ways to interpret this predictive distribution. Several works have explored the connections between Bayesian inference and deep ensembles (Ciosek et al., 2020; Gustafsson et al., 2020; He et al., 2020; Pearce et al., 2020; Wilson & Izmailov, 2020; Izmailov et al., 2021b; D’Angelo & Fortuin, 2021; D’Angelo et al., 2021). One interpretation views Equation 4 as a Monte Carlo approximation of Equation 5, where the posterior of the parameters is represented as a set of point masses centered at different modes, which may be viewed as approximate posterior samples. However, this interpretation is not the most insightful.
It is more enlightening to view approximate inference as the task of accurately approximating the integral in Equation 5. From this perspective, the focus is not on collecting posterior samples. For a fixed computational budget, a Monte Carlo average of predictive distribution values based on exact posterior samples can provide a poor approximation of the integral relative to alternatives. A more compelling approach to numerical integration is to choose parameter values that represent typical points in the posterior, indicative of regions with significant posterior probability mass, and that yield diverse predictions on the test set. A heuristic to achieve this goal is to choose points corresponding to different posterior modes, as achieved by deep ensembles (Wilson & Izmailov, 2020). In practice, there is more functional variability across different posterior modes compared to samples in the vicinity of a single mode, such as the ones found from a variational Gaussian approximation of the posterior.
@@ A.4 Posterior Sampling Algorithms
Simulating the following continuous-time stochastic differential equation (SDE) produces samples with $p(\boldsymbol{\theta}\mid\mathcal{D})$ as its stationary distribution:
@@
where $\alpha>0$ is the step size parameter and $\boldsymbol{\xi}$ is standard Gaussian noise. This discrete-time algorithm is known as the unadjusted Langevin algorithm, or the Langevin Monte Carlo algorithm. However, unlike the continuous-time Langevin diffusion equation, the discrete-time unadjusted Langevin algorithm does not simulate samples with $p(\boldsymbol{\theta}\mid\mathcal{D})$ as its stationary distribution, but instead produces samples that are only approximately drawn from $p(\boldsymbol{\theta}\mid\mathcal{D})$. The discretization of the SDE leads to a bias in the posterior samples, which can be reduced by decreasing the step size parameter $\alpha$.
@@
Other notable variants of SG-MCMC include stochastic gradient HMC (SG-HMC; Chen et al., 2014), which accelerates convergence using auxiliary momentum variables, and cyclical SG-MCMC (Zhang et al., 2020b), which employs a cyclical step size schedule to efficiently explore multiple modes of the parameter posterior distribution. There have also been efforts to mitigate the bias in SG-MCMC methods by using Metropolis adjustments (Zhang et al., 2020a; Garriga-Alonso & Fortuin, 2021).
@@ A.5 Prior Specification
which corresponds to $l_{2}$ regularization in the sense of maximum a posteriori (MAP) solutions with $\lambda=1/(2\sigma^{2})$. Thus, the larger the prior variance, the less regularization is incorporated, and vice versa. Combined with specific activation functions, such as the logistic function, which is close to linear around 0, choosing a small $\sigma^{2}$ results in more linear behavior of the neurons and their compositions, while large $\sigma^{2}$ allows for more non-linear behavior. Thus, the popular approach of choosing standard Gaussian priors is not satisfactory in most cases and may lead to misspecified models. This, in turn, can cause the cold posterior effect that has been known to be the case for linear models (Grünwald & Van Ommen, 2017), but is also observed for BNNs (Wenzel et al., 2020; Fortuin et al., 2022; Nabarro et al., 2022). For a specific problem, $\sigma^{2}$ can be chosen via hyper-parameter tuning or empirical Bayes. Moreover, a direct translation of the tuned $\sigma^{2}$ for some architecture (or, equivalently, $\lambda$ for frequentist neural networks) is possible. Another approach is to impose an inverse-Gamma hyper-prior on $\sigma^{2}$, for example, $\sigma^{2}\sim\Gamma^{-1}(\alpha,\beta)$, see Lampinen & Vehtari (2001). To incorporate prior dependencies between the parameters, i.i.d. Gaussian priors can be extended to multivariate normals with a zero mean vector and a covariance matrix $\boldsymbol{\Sigma}$, i.e.,
@@ A.7 Deep Kernel Machines
The objective (10) is maximized when $\mathbf{G}_{j}=\mathbf{K}(\mathbf{G}_{j-1})$, which is the same as the corresponding NNGP (Lee et al., 2017; Agrawal et al., 2020). If tempering is carried out according to the width with $\beta_{j}=1/r$, then the following objective is obtained:
@@
A model that optimizes the objective (11) is called a DKM, and (11) is known as the DKM objective. In the limit, the DKM objective does not depend on intermediate features $\mathbf{F}_{j}$, which means that learned representations in a DKM are described entirely by deterministic Gram matrices $\mathbf{G}_{1},\ldots,\mathbf{G}_{\eta}$. To interpret the DKM objective, notice that the likelihood term encourages data fitting, and the KL divergences regularize the model toward the NNGP (Lee et al., 2017; Agrawal et al., 2020). The amount of representation learning in the DKM can be controlled by varying the $\rho_{j}$ parameters. In contrast, the lack of likelihood term in the NNGP objective (10) prevents representation learning from occurring in the NNGP; intermediate Gram matrices are fixed and depend only on the input data.